Suppose a stock had an initial price of $70 per share, paid a dividend of $1.75 per share during the year, and had an ending share price of $77. Compute the percentage total return.

Multiple Choice

  • 12.50

  • 11.36

  • 14.00

  • 13.12
Answer #1

percentage total return=(Capital Gain +Dividend)/Initial investment

Capital gain =77-70=7

Dividend=1.75

percentage total return=(7+1.75)/70=12.50%